OpenStudy (anonymous):

SIG FIG QUESTION 1.5x10^3*534=801000 I round to two sig figs correct? And how would I do this with that number.

OpenStudy (whpalmer4):

\[1.5*10^3*534 = 1.5*10^3*5.34*10^2 = (1.5*5.34)*10^{3+2} = 8.01*10^5 \]rounding to 2 significant figures, you round the \(8.01\) to 2 significant figures, which is \(8.0\) giving you\[8.0*10^5\] If you wanted to write it without using scientific notation (where the number of significant figures in the mantissa indicates the number of significant figures for the whole thing), given that the least significant digit is a 0, you would have to use the notation where you draw a bar over the 0 digit which is the least significant digit, so it would look like\[8\bar{0}0000\]Personally, I hate that sort of notation.